OpenStudy (felicitynicolee):

If AB has endpoint A(2,13) and midpoint (10,-1) then give the coordinates of B.

OpenStudy (anonymous):

B(18, -15)

OpenStudy (felicitynicolee):

Can you tell me how i solve it though?

OpenStudy (anonymous):

Midpoint should have a coordinate value of mean of the two points

OpenStudy (anonymous):

10=(2+x)/2 20=2+x x=18 -1=(y+13)/2 -2=y+13 y=-15

OpenStudy (anonymous):

would that work?

OpenStudy (anonymous):

all jygerbs did is add the X1 and X2 and put it over 2 to find the midpoint. Since you don't know X2 you simply put X and solve for it. You then proceed to do the same for Y

OpenStudy (felicitynicolee):

By put it over your mean multiply?

OpenStudy (anonymous):

\[M(\frac{ x _{1}+x _{2} }{ 2 }, \frac{ y _{1}+y _{2} }{2 })\]

OpenStudy (anonymous):

I just substituted x1,x2,y1,y2 with the coordinates to solve the equation

OpenStudy (felicitynicolee):

Ohhhhh the midpoint formula. So you just plug it in. Okay thank you.

OpenStudy (anonymous):

award me a medal if you liked my response :)